# Mildly Accurate Computationally Differentially Private Inner Product Protocols Imply Oblivious Transfer

### v1

@@ 3.1 Distributions and Random Variables
The support of a distribution $P$ over a finite set ${\cal S}$ is defined by $\operatorname{Supp}(P):=\mathopen{}\mathclose{{}\left\{x\in{\cal S}:P(x)>0}\right\}$. For a distribution or a random variable $D$, let $d\leftarrow D$ denote that $d$ was sampled according to $D$. Similarly, for a set ${\cal S}$, let $x\leftarrow{\cal S}$ denote that $x$ is drawn uniformly from ${\cal S}$, and denote by ${\cal U}_{{\cal S}}$ the uniform distribution over ${\cal S}$. For a finite set ${\cal X}$ and a distribution $C_{X}$ over ${\cal X}$, we use the capital letter $X$ to denote the random variable that takes values in ${\cal X}$ and is sampled according to $C_{X}$. The statistical distance (also known as, variation distance) of two distributions $P$ and $Q$ over a discrete domain ${\cal X}$ is defined by $\mathsf{\textsc{SD}}\mathopen{}\mathclose{{}\left(P,Q}\right):=\max_{{\cal S}\subseteq{\cal X}}\mathopen{}\mathclose{{}\left|P({\cal S})-Q({\cal S})}\right|=\frac{1}{2}\sum_{x\in{\cal S}}\mathopen{}\mathclose{{}\left|P(x)-Q(x)}\right|$. For a vector $x=(x_{1},\ldots,x_{n})$ and index $i\in[n]$, we let $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n})$ and $x^{(i)}=(x_{1},\ldots,x_{i-1},-x_{i},x_{i+1},\ldots,x_{n})$, for a set ${\cal S}\subseteq[n]$ we let $x_{{\cal S}}=(x_{i})_{i\in{\cal S}}$ and $x_{-{\cal S}}=(x_{i})_{i\in[n]\setminus{\cal S}}$, and for a vector $r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ we let $x_{r}=(x_{i})_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=1}\right\}}$ and $x_{-r}=(x_{i})_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=0}\right\}}$.
@@ 3.2.1 Computational Differential Privacy
There are several ways for defining computational differential privacy (see Appendix A). We use the most relaxed version due to [BNO08]. For notational convenience, we focus on databases over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$.
@@ 3.3 Useful Claims
[…]
In this section, we state generic lemmas and propositions that we will use later in our proofs.
The following lemma which we prove in Section D.1, measures the distance between two uniform stings conditioned one a random index $i$ either being fixed to 00 or to $1$.
[…]
where the expectations are taken over $R$ and the randomness of $f$.
The following two propositions state that given the output of a differentially private function, it is not possible to predict well even a random index (even if all other indexes are leaked). The first proposition handles the information-theoretic case and the second handles the computation case. Both propositions are proven in Section D.2.
@@ 3.4.4 Approximate Weak Erasure Channel (AWEC)
We will make use of the following lemma, which shows that for some choices of the parameters, AWEC implies WEC. The lemma is proven in Appendix C.
@@ 4.1 Information-Theoretic Case
By Section 4.1 there exists a ppt protocol $\Gamma\coloneqq\Gamma^{C}$ such that channel $\tilde{C}\coloneqq C_{\Gamma(\kappa)}$ is a $(\ell,\alpha=0.001,p=0.001,q=0.001)$-AWEC. By Section 3.4.4 (and Section 3.3) there exists a ppt protocol $\Lambda\coloneqq\Lambda^{\tilde{C}}$ such that the channel $\widehat{C}=C_{\Lambda(\kappa)}$ is a $(\alpha^{\prime}=0.002,\>p^{\prime}=0.001,\>q^{\prime}=1/2+0.022))$-WEC. Since $44(\varepsilon^{\prime}+p^{\prime})<1-q^{\prime}$, by Theorem 3.1 there exists a ppt protocol $\Pi$, such that $\Pi^{\widehat{C}}$ is a semi-honest statistically secure OT, concluding the proof.     $\Box$
@@ 4.2 Computational Case
Proof of 4.2.0.1.
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$. By 4.2.0.1, $\widehat{\Psi}\coloneqq\widehat{\Psi}^{\Psi}$ induces a channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{\kappa}=((X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa}))}\right\}_{\kappa\in{\mathbb{N}}}$ that is $(\varepsilon,\delta)$-CDP and $(\ell,0.999)$-acccurate for the inner-product functionality. Let $\Gamma$ be the ppt protocol guarrented by Section 4.1, we now claim that $\Gamma^{\Psi}(\kappa)\coloneqq\Gamma^{C_{\kappa}}(\kappa)$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. Assume towards contradiction that this does not hold, then by Section 4.1, there exists a ppt $\mathsf{D}$, such that for infinity many $\kappa\in{\mathbb{N}}$, $\mathsf{D}$ contradicts one of the two secrecy properties of ${\sf CompAWEC}$. Fix such $\kappa\in{\mathbb{N}}$ and omit it from the notation when clear from the context, and without loss of generality, assume that $\mathsf{D}$ contradicts the first secrecy property of ${\sf CompAWEC}$ (the case where $D$ contradicts the second property is essentially identical). By the first item of Section 4.1, there exists a ppt algorithm $\tilde{\mathsf{A}}$ such that $Y^{*}_{i}={\tilde{\mathsf{A}}}^{\mathsf{D}}(i,\>Y_{-i},\>X,\>U)$ and it holds that:
@@
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality. By 4.2.0.1 there exists a ppt protocol $\Gamma$ such that $\Gamma\coloneqq\Gamma^{\Psi}$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. By Section 3.4.4, there exists a ppt protocol $\Lambda$ such that $\Lambda\coloneqq\Lambda^{\Gamma}$ is $(\epsilon^{\prime}=\epsilon+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2(q+0.01))$-CompWEC. Finally, since $44(\varepsilon^{\prime}+p^{\prime})<1-q^{\prime}$ by Theorem 3.2 there exists a ppt protocol $\Pi$, such that $\Pi^{\Lambda}$ is a semi-honest computationally secure OT, as required.     $\Box$
@@
We next use the following lemma, proven in Section D.3 (recall that a proof overview appears in Section 2.1).
@@
We now ready to finalize the proof of Property 1 of Section 4.1 using Equations 13 and 5.1.
@@
By Equations 16 and 17, assuming that $\delta\leq 1/n$ where $n$ is large enough, there exists a constant $c>0$ such that if $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$ then ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Y_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}\neq Y_{i}}\right]}\right]+\delta$, as required. Recall that $k=\mathopen{}\mathclose{{}\left\lceil e^{\lambda_{1}\varepsilon}\lambda_{2}\cdot\ell^{2}}\right\rceil$, where $\lambda_{1},\lambda_{2}$ are the constants from Equation 29. Hence, we can set a bound of $e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$ on $\ell$ for $c_{1}=\lambda_{1}/2+1/6$ and $c_{2}=\sqrt{c/\lambda_{2}}$ to guarantee that $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$.
@@
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section D.3 with the function $f_{z,w}(r)=f(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
@@
We next prove Equations 36 and 37 using Equations 39 and 40.
@@
We next prove the lemma using Equations 41 and 42.
@@
It is left to prove Equation 37. Observe that Equations 41 and 42 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that
@@
We deduce from Sections 5.2 and 28 that
@@
Appendix A Related Works
Appendix B [HaitnerMST22]’s Protocol Cannot Imply ${\sf OT}$
In this section, we show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in [HaitnerMST22]’s protocol (Appendix B) can be simulated using a trivial protocol, without using $\Pi$ at all.
@@ Appendix A Related Works
Oracle: An accurate ${\sf CDP}$ protocol $\Pi$ for the inner-product. Operation: 1. $\mathsf{A}$ and $\mathsf{B}$ choose random inputs $x\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$, respectively. 2. The parties interact using $\Pi$ to get approximation $z$ of $\langle x,y\rangle$. 3. $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=0}\right\}$. 4. Finally, $\mathsf{A}$ computes and outputs ${\rm out}_{\mathsf{A}}=\langle x_{-r},y_{-r}\rangle$, and $\mathsf{B}$ outputs ${\rm out}_{\mathsf{B}}=z-\langle x_{r},y_{r}\rangle$.
To see this, assume that $\Pi$ is a protocol that on inputs $x$ and $y$, outputs $z=\langle x,y\rangle+e_{\mathsf{A}}+e_{\mathsf{B}}$, where $e_{\mathsf{A}}$ and $e_{\mathsf{B}}$ are independent samples from the ${\rm Lap}(2/\varepsilon)$ distribution. Moreover, assume that $\Pi$ reveals $e_{\mathsf{A}}$ to $\mathsf{A}$ and $e_{\mathsf{B}}$ to $\mathsf{B}$ (and nothing else). Such a protocol is indeed differential private, and it can be implemented using secure multi-party computation. Moreover, as we show in this work, such a protocol can be used to construct OT. However, when composed with the ${\sf KA}$ protocol of [HaitnerMST22], the resulting protocol can be simulated trivially. [^14] More formally, and using the definition given in Section 3.4, we claim that the channel induced by executing Appendix B with oracle access to the channel $\Pi=\mathopen{}\mathclose{{}\left\{((x,(\langle x,y\rangle+e_{A}+e_{B},e_{A})),(y,(\langle x,y\rangle+e_{A}+e_{B},e_{B})))}\right\}_{x,y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n},e_{A},e_{B}\leftarrow{\rm Lap}(2/\epsilon)}$ is a trivial channel.
@@
Clearly, Appendix B is a trivial protocol and does not use any cryptographic assumptions. However, the views of $\mathsf{A}$ and $\mathsf{B}$ in $\Pi^{\prime}$ contain all the information learned by the parties in the ${\sf KA}$ protocol ($(x,z,e_{A},r,y_{-r})$ and $(y,z,e_{B},r,x_{r})$ respectively). Moreover, we claim that the parties in $\Pi^{\prime}$ do not learn any information that the parties in the ${\sf KA}$ protocol did not learn. Indeed, the only new value learned by $\mathsf{B}$, $z_{A},$ can be also computed by $\mathsf{B}$ in the ${\sf KA}$ protocol by computing $z-e_{B}-\langle x_{r},y_{r}\rangle$. Similarly, $z_{B}$ can be already computed by $\mathsf{A}$.
We note that every protocol that uses only a communication channel cannot be used to construct ${\sf OT}$ unless ${\sf OT}$ already exists, and similarly, every protocol that uses $\Pi^{\prime}$ as a subroutine (in a black-box manner) cannot be used to construct ${\sf OT}$. Since Appendix B (that is, the views and outputs of the parties when running Appendix B) are the same as $\Pi^{\prime}$, we conclude that Appendix B could not be used to construct ${\sf OT}$ in a black-box manner.
Appendix C WEC from AWEC (Proof of Section 3.4.4)

### v2

@@ 4.1 Distributions and Random Variables
The support of a distribution $P$ over a finite set ${\cal S}$ is defined by $\operatorname{Supp}(P):=\mathopen{}\mathclose{{}\left\{x\in{\cal S}:P(x)>0}\right\}$. For a distribution or a random variable $D$, let $d\leftarrow D$ denote that $d$ was sampled according to $D$. Similarly, for a set ${\cal S}$, let $x\leftarrow{\cal S}$ denote that $x$ is drawn uniformly from ${\cal S}$, and denote by ${\cal U}_{{\cal S}}$ the uniform distribution over ${\cal S}$. For a finite set ${\cal X}$ and a distribution $C_{X}$ over ${\cal X}$, we use the capital letter $X$ to denote the random variable that takes values in ${\cal X}$ and is sampled according to $C_{X}$. The statistical distance (also known as, variation distance) of two distributions $P$ and $Q$ over a discrete domain ${\cal X}$ is defined by $\mathsf{\textsc{SD}}\mathopen{}\mathclose{{}\left(P,Q}\right):=\max_{{\cal S}\subseteq{\cal X}}\mathopen{}\mathclose{{}\left|P({\cal S})-Q({\cal S})}\right|=\frac{1}{2}\sum_{x\in{\cal S}}\mathopen{}\mathclose{{}\left|P(x)-Q(x)}\right|$. For a vector $x=(x_{1},\ldots,x_{n})$ and index $i\in[n]$, we let $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n})$ and $x^{(i)}=(x_{1},\ldots,x_{i-1},-x_{i},x_{i+1},\ldots,x_{n})$, for a set ${\cal S}\subseteq[n]$ we let $x_{{\cal S}}=(x_{i})_{i\in{\cal S}}$ and $x_{-{\cal S}}=(x_{i})_{i\in[n]\setminus{\cal S}}$, and for a vector $r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ we let $x_{r}=x_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=1}\right\}}$ and $x_{-r}=x_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=0}\right\}}$.
@@ 4.2.1 Computational Differential Privacy
There are several ways for defining computational differential privacy (see Section 3). We use the most relaxed version due to [BNO08]. For notational convenience, we focus on databases over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$.
@@ 4.3 Useful Claims
The following two propositions state that given the output of a differentially private function, it is not possible to predict well even a random index (even if all other indexes are leaked). The first proposition handles the information-theoretic case and the second handles the computation case. Both propositions are proven in Section C.1.
@@ 4.4.4 Approximate Weak Erasure Channel (AWEC)
We will make use of the following lemma, which shows that for some choices of the parameters, AWEC implies WEC. The lemma is proven in Appendix B.
@@ 5.1 Information-Theoretic Case
By Section 5.1 there exists a ppt protocol $\Gamma\coloneqq\Gamma^{C}$ such that channel $\tilde{C}\coloneqq C_{\Gamma(\kappa)}$ is a $(\ell,\alpha=0.001,p=0.001,q=0.001)$-AWEC. By Section 4.4.4 (and Section 4.3) there exists a ppt protocol $\Lambda\coloneqq\Lambda^{\tilde{C}}$ such that the channel $\widehat{C}=C_{\Lambda(\kappa)}$ is a $(\alpha^{\prime}=0.002,\>p^{\prime}=0.001,\>q^{\prime}=1/2+0.022))$-WEC. Since $44(\alpha^{\prime}+p^{\prime})<1-q^{\prime}$, by Theorem 4.1 there exists a ppt protocol $\Pi$, such that $\Pi^{\widehat{C}}$ is a semi-honest statistically secure OT, concluding the proof.     $\Box$
@@ 5.2 Computational Case
Proof of 5.2.0.1.
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$. By 5.2.0.1, $\widehat{\Psi}\coloneqq\widehat{\Psi}^{\Psi}$ induces a channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{\kappa}=((X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa}))}\right\}_{\kappa\in{\mathbb{N}}}$ that is $(\varepsilon,\delta)$-CDP and $(\ell,0.999)$-acccurate for the inner-product functionality. Let $\Gamma$ be the ppt protocol guarrented by Section 5.1, we now claim that $\Gamma^{\Psi}(\kappa)\coloneqq\Gamma^{C_{\kappa}}(\kappa)$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. Assume towards contradiction that this does not hold, then by Section 5.1, there exists a ppt $\mathsf{D}$, such that for infinity many $\kappa\in{\mathbb{N}}$, $\mathsf{D}$ contradicts one of the two secrecy properties of ${\sf CompAWEC}$. Fix such $\kappa\in{\mathbb{N}}$ and omit it from the notation when clear from the context, and without loss of generality, assume that $\mathsf{D}$ contradicts the first secrecy property of ${\sf CompAWEC}$ (the case where $D$ contradicts the second property is essentially identical). By the first item of Section 5.1, there exists a ppt algorithm $\tilde{\mathsf{A}}$ such that $Y^{*}_{i}={\tilde{\mathsf{A}}}^{\mathsf{D}}(i,\>Y_{-i},\>X,\>U)$ and it holds that:
@@
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality. By 5.2.0.1 there exists a ppt protocol $\Gamma$ such that $\Gamma\coloneqq\Gamma^{\Psi}$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. By Section 4.4.4, there exists a ppt protocol $\Lambda$ such that $\Lambda\coloneqq\Lambda^{\Gamma}$ is $(\epsilon^{\prime}=\epsilon+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2(q+0.01))$-CompWEC. Finally, since $44(\varepsilon^{\prime}+p^{\prime})<1-q^{\prime}$ by Theorem 4.2 there exists a ppt protocol $\Pi$, such that $\Pi^{\Lambda}$ is a semi-honest computationally secure OT, as required.     $\Box$
@@
We next use the following lemma, proven in Section C.2 (recall that a proof overview appears in Section 2.3).
@@
We now ready to finalize the proof of Property 1 of Section 5.1 using Equations 12 and 6.1.
@@
By Equations 15 and 16, assuming that $\delta\leq 1/n$ where $n$ is large enough, there exists a constant $c>0$ such that if $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$ then ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Y_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}\neq Y_{i}}\right]}\right]+\delta$, as required. Recall that $k=\mathopen{}\mathclose{{}\left\lceil e^{\lambda_{1}\varepsilon}\lambda_{2}\cdot\ell^{2}}\right\rceil$, where $\lambda_{1},\lambda_{2}$ are the constants from Equation 28. Hence, we can set a bound of $e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$ on $\ell$ for $c_{1}=\lambda_{1}/2+1/6$ and $c_{2}=\sqrt{c/\lambda_{2}}$ to guarantee that $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$.
@@
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section C.2 with the function $f_{z,w}(r)=f(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
@@
We next prove Equations 34 and 35 using Equations 37 and 38.
@@
We next prove the lemma using Equations 39 and 40.
@@
It is left to prove Equation 35. Observe that Equations 39 and 40 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that
@@
We deduce from Sections 6.2 and 27 that
@@
Appendix A [HaitnerMST22]’s Protocol Cannot Imply ${\sf OT}$
In this section, we show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in [HaitnerMST22]’s protocol (Appendix A) can be simulated using a trivial protocol, without using $\Pi$ at all.
@@
Oracle: An accurate ${\sf CDP}$ protocol $\Pi$ for the inner-product. Operation: 1. $\mathsf{A}$ and $\mathsf{B}$ choose random inputs $x\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$, respectively. 2. The parties interact using $\Pi$ to get approximation $z$ of $\langle x,y\rangle$. 3. $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$. 4. Finally, $\mathsf{A}$ computes and outputs ${\rm out}_{\mathsf{A}}=\langle x_{-r},y_{-r}\rangle$, and $\mathsf{B}$ outputs ${\rm out}_{\mathsf{B}}=z-\langle x_{r},y_{r}\rangle$.
To see this, assume that $\Pi$ is a protocol that on inputs $x$ and $y$, outputs $z=\langle x,y\rangle+e_{\mathsf{A}}+e_{\mathsf{B}}$, where $e_{\mathsf{A}}$ and $e_{\mathsf{B}}$ are independent samples from the ${\rm Lap}(2/\varepsilon)$ distribution. Moreover, assume that $\Pi$ reveals $e_{\mathsf{A}}$ to $\mathsf{A}$ and $e_{\mathsf{B}}$ to $\mathsf{B}$ (and nothing else). Such a protocol is indeed differential private, and it can be implemented using secure multi-party computation. Moreover, as we show in this work, such a protocol can be used to construct OT. However, when composed with the ${\sf KA}$ protocol of [HaitnerMST22], the resulting protocol can be simulated trivially. [^14] More formally, and using the definition given in Section 4.4, we claim that the channel induced by executing Appendix A with oracle access to the channel $\Pi=\mathopen{}\mathclose{{}\left\{((x,(\langle x,y\rangle+e_{A}+e_{B},e_{A})),(y,(\langle x,y\rangle+e_{A}+e_{B},e_{B})))}\right\}_{x,y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n},e_{A},e_{B}\leftarrow{\rm Lap}(2/\epsilon)}$ is a trivial channel.
@@
Clearly, Appendix A is a trivial protocol and does not use any cryptographic assumptions. However, the views of $\mathsf{A}$ and $\mathsf{B}$ in $\Pi^{\prime}$ contain all the information learned by the parties in the ${\sf KA}$ protocol ($(x,z,e_{A},r,y_{-r})$ and $(y,z,e_{B},r,x_{r})$ respectively). Moreover, we claim that the parties in $\Pi^{\prime}$ do not learn any information that the parties in the ${\sf KA}$ protocol did not learn. Indeed, the only new value learned by $\mathsf{B}$, $z_{A},$ can be also computed by $\mathsf{B}$ in the ${\sf KA}$ protocol by computing $z-e_{B}-\langle x_{r},y_{r}\rangle$. Similarly, $z_{B}$ can be already computed by $\mathsf{A}$.
We note that every protocol that uses only a communication channel cannot be used to construct ${\sf OT}$ unless ${\sf OT}$ already exists, and similarly, every protocol that uses $\Pi^{\prime}$ as a subroutine (in a black-box manner) cannot be used to construct ${\sf OT}$. Since Appendix A (that is, the views and outputs of the parties when running Appendix A) are the same as $\Pi^{\prime}$, we conclude that Appendix A could not be used to construct ${\sf OT}$ in a black-box manner.
Appendix B WEC from AWEC (Proof of Section 4.4.4)